in a certain company 140 of the employees are men. what is the total number of employees if 5 out of every 7 employees are men
Answer by jim_thompson5910(35256)

5/7 = 140/x

5x = 7*140

5x = 980

x = 980/5

x = 196

So there are 196 employees total.